The mean NAT scores of Grade 10 students is 65. Sixty (60) students were chosen and

found that the standard deviation of their scores is 5. What is the probability that their

mean is between 64 and 67?


1
Expert's answer
2022-04-25T17:52:36-0400

"P(64<X<67)=P(\frac{64-65}{5/\sqrt{60}}<Z<\frac{67-65}{5/\sqrt{60}})=P(-1.55<Z<3.10)=0.999-0.0606=0.9384"